# Dyson Equation for Correlated Linearizations and Test Error of Random Features Regression

### v2

@@ 2.4.5 Gaussian Equivalence
Theorem 2.4 establishes a Gaussian equivalence principle, indicating that every random features model trained with ridge regression, as described in the statement of Theorem 2.4, performs equivalently to a surrogate Gaussian model with a matching covariance structure. However, it is important to note, as mentioned in [46], that the distribution of the input data can impact the performance of the random features model. This influence stems from the fact that, although there is Gaussian equivalence at the level of random feature matrices, the distribution of the input may influence the covariance matrices $K_{AA^{T}}$, $K_{A\hat{A}^{T}}$, $K_{\hat{A}A^{T}}$, and $K_{\hat{A}\hat{A}^{T}}$, which are directly linked to the performance of the random features model.
@@ 2.4.6 Global Anisotropic Law for Rectangular Random Matrices
$L^{-1}=\begin{bmatrix}R&RA&RA\hat{A}^{T}&0\\ A^{T}R&\bar{R}&\bar{R}\hat{A}^{T}&0\\ \hat{A}A^{T}R&\hat{A}\bar{R}&\hat{A}\bar{R}\hat{A}^{T}&-I_{n_{\text{test}}}\\ 0&0&-I_{n_{\text{test}}}&0\end{bmatrix},$
where $R=(AA^{T}+\delta I_{n_{\text{train}}})^{-1}$, $\bar{R}=-\delta(\delta I_{d}+A^{T}A)^{-1}$, and $M$ is given in (23).
For any $(i,j)\in\{1,2,3,4\}^{2}$, consider $U$ as a block matrix with $U_{i,j}$ in the $(j,i)$th block and zeros elsewhere. This construction leads to $\mathrm{tr}(U(L^{-1}-M(0)))=U_{i,j}(L^{-1})_{i,j}-U_{i,j}M(0)_{j,i}\to 0$ almost surely as $n\to\infty$. Moreover, the matrix $U^{T}U$ is block diagonal, with $U_{i,j}^{T}U_{i,j}$ in the $(i,i)$-th block and zeros elsewhere, ensuring that $\|U\|_{\ast}\leq 1$ whenever $\|U_{i,j}\|_{\ast}\leq 1$. This implies that $\mathrm{tr}(U((L^{-1})_{i,j}-M_{i,j}(0)))\to 0$ almost surely as $n\to\infty$, for any sequence $U$ of the same size as $(L^{-1})_{i,j}$ and $M_{i,j}(0)$, with $\|U\|_{\ast}\leq 1$. This result constitutes a global anisotropic law for every block of $L^{-1}$, providing deterministic equivalents for products of $A$, $\hat{A}$, and related resolvents.
@@ 2.4.7 Implicit Regularization and Relation to Kernel Regression
Theorem 2.4 demonstrates the concentration of the empirical test error of random features ridge regression around the deterministic quantity $d\beta\|K_{AA^{T}}^{1/2}My\|^{2}+\|d\alpha K_{\hat{A}A^{T}}My+\hat{y}\|^{2}$ as the dimensions increases in the proportional regime. The second term, $\|d\alpha K_{\hat{A}A^{T}}My+\hat{y}\|^{2}=\|\hat{y}-dK_{\hat{A}A^{T}}(dK_{AA^{T}}+(-\delta/\alpha)I_{n_{\text{train}}})^{-1}y\|^{2}$, is equivalent to the squared norm of the empirical test error for kernel ridge regression with ridge parameter $-\delta/\alpha\in\mathbb{R}_{>0}$ and the conjugate kernel $K(x_{1},x_{2})=\frac{d}{n}\mathbb{E}_{z\sim\mathcal{N}(0,I_{n_{0}})}[\sigma(x_{1}^{T}\varphi(z))\sigma(x^{T}_{2}\varphi(z))]$. We will show in Section 6 that $-1\leq\alpha<0$, which implies that $\delta<-\delta/\alpha$. This reveals that the randomness in the random features matrix acts as a form of regularization, similar to increasing the ridge parameter in kernel ridge regression. This is reminiscent of [42] and the generalization in [20], and is related to the implicit regularization of the random features model. In fact, the proof of Theorem 2.4 recovers both of those results.
However, the additional term $d\beta\|K_{AA^{T}}^{1/2}My\|^{2}$ represents the variance in the empirical test error due to the random weights. Consequently, despite the computational benefits of the random features ridge regression compared to kernel ridge regression, Theorem 2.4 indicates that in the proportional regime, the random features model may still underperform kernel ridge regression if $d\beta\|K_{AA^{T}}^{1/2}My\|^{2}$ is significant.
@@ 2.4.8 Numerical Considerations
Theorem 2.4 suggests that computing the asymptotic deterministic equivalent for $E_{\text{test}}$ can be reduced to solving a scalar fixed-point equation. As shown in Lemma 6.10, the iterates $\{\alpha_{k}\}_{k\in\mathbb{N}_{0}}$ obtained by iterating $\alpha_{k+1}=-(1+\mathrm{tr}(K_{AA^{T}}(\delta I_{n_{\text{train}}}-\alpha_{k}dK_{AA^{T}})^{-1}))^{-1}$ for every $k\in\mathbb{N}$ with arbitrary $\alpha_{0}\in\mathbb{R}_{\leq 0}$ converge to $\alpha$ as $k\to\infty$. Using the spectral decomposition $K_{AA^{T}}=U\mathrm{diag}\{\lambda_{j}(K_{AA^{T}})\}_{j=1}^{n_{\text{train}}}U^{T}$ for some orthonormal matrix $U$, we can rewrite the iteration as
$\alpha_{k+1}=-\left(1+\sum_{j=1}^{n_{\text{train}}}\frac{\lambda_{j}(K_{AA^{T}})}{\delta-d\alpha_{k}\lambda_{j}(K_{AA^{T}})}\right)^{-1}.$
Hence, instead of performing a matrix inversion at each iteration, we compute one spectral decomposition and use the above formula to update $\alpha_{k}$ for every $k\in\mathbb{N}$. This allows us to more efficiently compute the deterministic equivalent of $E_{\text{test}}$ for various activation functions and dimensions. Moreover, when $\varphi$ is the identity, the kernel matrices $K_{AA^{T}}$, $K_{A\hat{A}^{T}}$, and $K_{\hat{A}\hat{A}^{T}}$ can be computed without Monte Carlo simulations using the closed-form expressions in [46, Table 1].
@@ 3.1 Conjugate Kernel
The conjugate kernel, defined as the Gram matrix of features produced by the final layer of a network, is central to the analysis of random features models [27]. This connection stems from the fact that the ouput of a neural network is linear in those derived features. Thus, the conjugate kernel characterizes the training and test error of this linear model. In the case of shallow random features models, the conjugate kernel is equivalent to the Gram matrix of the random features themselves.
Numerous studies have employed random matrix theory to investigate the conjugate kernel in the proportional regime. Works such as [53, 17] leverage the moment method to establish deterministic equivalents for random features models with isotropic data and weight matrices. [54] extend these results to include an additive bias. To address more realistic data distributions, [27] study the case of nearly orthogonal data. They introduce a notion of orthogonality that can propagate through network layers, providing control over key quantities related to the conjugate kernel. Notably, their settings encompass the case of isotropic data, with the difference that the author conditioned on the data. The conjugate kernel is also studied in [20, 46] utilizing concentration of measure and leave-one-out techniques. Beyond bulk spectrum analysis, motivated in part by results from [14, 21], where a single gradient step in a neural network at random initialization is shown to be equivalent to a spiked random features model, recent works investigate outlier eigenvalues of the conjugate kernel  [18, 61]. The structure in random features model turns out to have important implication, notably in understanding the mechanisms of feature learning.
The proof of our main result recovers a deterministic equivalent for the conjugate kernel, consistent with the findings of [20, 46], employing a significantly different approach. In particular, replacing the ridge parameter by a spectral parameter in Theorem 2.4 and using the Stieltjes inversion lemma, we can recover the density of the empirical spectral distribution of the conjugate kernel.
@@ 3.2 Dyson Equation and Linearizations
The Dyson equation is a self-consistent equation that has proven to be a valuable tool in random matrix theory. For an excellent introduction to this subject, we recommend [24]. The vector version of the Dyson equation has proven invaluable in establishing local laws for Generalized Wigner matrices [4, 6, 7]. Extending its applicability, the matrix Dyson equation has been instrumental in investigating local laws for Hermitian matrices with correlations featuring fast decay, as well as those with slower correlation decay, particularly focusing on regular edges [11] and regions away from the support edges [25]. A notable advantage of the Dyson equation is that its solution is given by a Stieltjes representation. Leveraging this equation, detailed regularity properties of the self-consistent density of states have been explored [9], and bounds on the spectrum of Kronecker random matrices have been established [10].
The Dyson equation is a useful tool for deriving deterministic equivalents, but it is limited to certain classes of “Wigner-like” matrices. In order to extend its applicability to a broader class of matrices, we employ a linearization trick. The concept of linearization became particularly important following the influential work of [33]. This work essentially demonstrated that to analyze a polynomial expression in matrices, it is sufficient to consider a linear polynomial with matrix coefficients. However, a limitation of [33] is that their linearization approach does not inherently preserve the self-adjointness of the original polynomial expression. [12] addressed this problem, proving that a self-adjoint polynomial expression allows the linearization’s coefficients to be chosen in a way that retains self-adjointness [12]. This finding was later generalized to include rational expressions [38]. The fact that linearizations are not unique has inspired other linearization methods. For example, [26] developed the concept of a minimal linearization. A key strength of these linearization techniques is that their supporting arguments are often constructive, giving explicit instructions for creating appropriate linearizations.
The linearization trick is a powerful tool within free probability, enabling the study of random matrix polynomials on the global scale [12, 15, 33, 38, 39] and the local scale [26, 28, 13]. Combined with operator-valued free probability, the linearization trick—referred to as the pencil method in this context—has found successful applications in the study of simple neural networks [52, 1, 2, 3, 59].
Given a linearization $L\in\mathbb{C}^{\ell\times\ell}$ for $\ell=n+d$, we aim to understand the asymptotic properties of the pseudo-resolvent $(L-z\Lambda_{\ell})^{-1}$, where $\Lambda=\mathrm{diag}\{I_{n},0_{d\times d}\}$. The lack of a spectral parameter spanning the entire diagonal makes pseudo-resolvents more challenging to analyze compared to resolvents. For example, unlike Hermitian matrix resolvents which are bounded by the inverse of the imaginary part of the spectral parameter, pseudo-resolvents lack such an a priori bound. Despite these challenges, the Dyson equation has been extended for pseudo-resolvent analysis. [12] derived global laws for linearizations of polynomials in independent Wigner matrices, terming their Dyson equation the Schwinger-Dyson equation [12]. This work was expanded to study the anticommutator on a local scale and more generally to polynomials of matrices with independent centered entries and suitable normalization under the name Dyson equation for linearizations (DEL) [13, 26, 28].

### v4

@@ 2.4.5 Gaussian Equivalence
Theorem 2.3 establishes a Gaussian equivalence principle, indicating that every random features model trained with ridge regression, as described in the statement of Theorem 2.3, performs equivalently to a surrogate Gaussian model with a matching covariance structure. However, it is important to note, as mentioned in [louart_random-matrix-approach], that the distribution of the input data can impact the performance of the random features model. This influence stems from the fact that, although there is Gaussian equivalence at the level of random feature matrices, the distribution of the input may influence the covariance matrices $K_{AA^{\top}}$, $K_{A\hat{A}^{\top}}$, $K_{\hat{A}A^{\top}}$, and $K_{\hat{A}\hat{A}^{\top}}$, which are directly linked to the performance of the random features model.
@@ 2.4.6 Global Anisotropic Law for Rectangular Random Matrices
$L^{-1}=\begin{bmatrix}R&RA&RA\hat{A}^{\top}&0\\ A^{\top}R&\bar{R}&\bar{R}\hat{A}^{\top}&0\\ \hat{A}A^{\top}R&\hat{A}\bar{R}&\hat{A}\bar{R}\hat{A}^{\top}&-I_{n_{\text{test}}}\\ 0&0&-I_{n_{\text{test}}}&0\end{bmatrix}$,
where $R=(AA^{\top}+\delta I_{n_{\text{train}}})^{-1}$, $\bar{R}=-\delta(\delta I_{d}+A^{\top}A)^{-1}$, and $M$ is given in (23).
For any $(i,j)\in\{1,2,3,4\}^{2}$, consider $U$ as a block matrix with $U_{i,j}$ in the $(j,i)$th block and zeros elsewhere. This construction leads to $\mathrm{tr}(U(L^{-1}-M(0)))=U_{i,j}(L^{-1})_{i,j}-U_{i,j}M(0)_{j,i}\to 0$ almost surely as $n\to\infty$. Moreover, the matrix $U^{\top}U$ is block diagonal, with $U_{i,j}^{\top}U_{i,j}$ in the $(i,i)$-th block and zeros elsewhere, ensuring that $\|U\|_{\ast}\leq 1$ whenever $\|U_{i,j}\|_{\ast}\leq 1$. This implies that $\mathrm{tr}(U((L^{-1})_{i,j}-M_{i,j}(0)))\to 0$ almost surely as $n\to\infty$, for any sequence $U$ of the same size as $(L^{-1})_{i,j}$ and $M_{i,j}(0)$, with $\|U\|_{\ast}\leq 1$. This result constitutes a global anisotropic law for every block of $L^{-1}$, providing deterministic equivalents for products of $A$, $\hat{A}$, and related resolvents.
@@ 2.4.7 Implicit Regularization and Relation to Kernel Regression
Theorem 2.3 demonstrates the concentration of the empirical test error of random features ridge regression around the deterministic quantity $d\beta\|K_{AA^{\top}}^{1/2}My\|^{2}+\|d\alpha K_{\hat{A}A^{\top}}My+\hat{y}\|^{2}$ as the dimensions increases in the proportional regime. The second term, $\|d\alpha K_{\hat{A}A^{\top}}My+\hat{y}\|^{2}=\|\hat{y}-dK_{\hat{A}A^{\top}}(dK_{AA^{\top}}+(-\delta/\alpha)I_{n_{\text{train}}})^{-1}y\|^{2}$, is equivalent to the squared norm of the empirical test error for kernel ridge regression with ridge parameter $\delta/\alpha\in\mathbb{R}_{>0}$ and the conjugate kernel $K(x_{1},x_{2})=\frac{d}{n}\mathbb{E}_{z\sim\mathcal{N}(0,I_{n_{0}})}[\sigma(x_{1}^{\top}\varphi(z))\sigma(x^{\top}_{2}\varphi(z))]$. We will show in Section 6 that $1\leq\alpha<0$, which implies that $\delta<-\delta/\alpha$. This reveals that the randomness in the random features matrix acts as a form of regularization, similar to increasing the ridge parameter in kernel ridge regression. This is reminiscent of [jacot_RF] and the generalization in [chouard2022quantitative], and is related to the implicit regularization of the random features model. In fact, the proof of Theorem 2.3 recovers both of those results.
However, the additional term $d\beta\|K_{AA^{\top}}^{1/2}My\|^{2}$ represents the variance in the empirical test error due to the random weights. Consequently, despite the computational benefits of the random features ridge regression compared to kernel ridge regression, Theorem 2.3 indicates that in the proportional regime, the random features model may still underperform kernel ridge regression if $d\beta\|K_{AA^{\top}}^{1/2}My\|^{2}$ is significant.
@@ 2.4.8 Numerical Considerations
Theorem 2.3 suggests that computing the asymptotic deterministic equivalent for $E_{\text{test}}$ can be reduced to solving a scalar fixed-point equation. As shown in Lemma 6.9, the iterates $\{\alpha_{k}\}_{k\in\mathbb{N}_{0}}$ obtained by iterating $\alpha_{k+1}=-(1+\mathrm{tr}(K_{AA^{\top}}(\delta I_{n_{\text{train}}}-\alpha_{k}dK_{AA^{\top}})^{-1}))^{-1}$ for every $k\in\mathbb{N}$ with arbitrary $\alpha_{0}\in\mathbb{R}_{\leq 0}$ converge to $\alpha$ as $k\to\infty$. Using the spectral decomposition $K_{AA^{\top}}=U\mathrm{diag}\{\lambda_{j}(K_{AA^{\top}})\}_{j=1}^{n_{\text{train}}}U^{\top}$ for some orthonormal matrix $U$, we can rewrite the iteration as
$\alpha_{k+1}=-\left(1+\sum_{j=1}^{n_{\text{train}}}\frac{\lambda_{j}(K_{AA^{\top}})}{\delta-d\alpha_{k}\lambda_{j}(K_{AA^{\top}})}\right)^{-1}.$
Hence, instead of performing a matrix inversion at each iteration, we compute one spectral decomposition and use the above formula to update $\alpha_{k}$ for every $k\in\mathbb{N}$. This allows us to more efficiently compute the deterministic equivalent of $E_{\text{test}}$ for various activation functions and dimensions. Moreover, when $\varphi$ is the identity, the kernel matrices $K_{AA^{\top}}$, $K_{A\hat{A}^{\top}}$, and $K_{\hat{A}\hat{A}^{\top}}$ can be computed without Monte Carlo simulations using the closed-form expressions in [louart_random-matrix-approach, Table 1].
@@ 3.1 Conjugate Kernel
Consider a fully-connected feedforward neural network with $L$ layers, input dimension $n_{0}=d$, hidden layer dimensions $\{n_{\ell}\}_{\ell=1}^{L-1}$, scalar output and activation functions $\sigma$. The network is defined as the parametric function
$f:x\in\mathbb{R}^{n_{0}}\mapsto w^{\top}\frac{1}{\sqrt{n_{L}}}\sigma\left(\frac{1}{\sqrt{n_{L-1}}}\sigma\left(\cdots\frac{1}{\sqrt{n_{1}}}\sigma\left(\frac{1}{\sqrt{n_{0}}}x^{\top}W_{1}\right)W_{2}\right)\cdots W_{L-1}\right)\in\mathbb{R},$
where $W_{\ell}\in\mathbb{R}^{n_{\ell-1}\times n_{\ell}}$ for every $\ell\in[L-1]$ and $w\in\mathbb{R}^{n_{L}}$ are the weights of the network and the activation function $\sigma$ is applied entrywise. Given $n$ datapoints $\{x_{i}\}_{i=1}^{n}$, we arrange them in a data matrix $X\in\mathbb{R}^{n\times n_{0}}$ and let $f(X)=[f(x_{1})^{\top},\ldots,f(x_{n})^{\top}]^{\top}\in\mathbb{R}^{n}$ be the vector of network outputs. Recursively define the matrices of post-activation by
$X_{0}=X,\quad X_{\ell}=\frac{1}{\sqrt{n_{\ell}}}\sigma\left(\frac{1}{\sqrt{n_{\ell-1}}}X_{\ell-1}W_{\ell}\right)\in\mathbb{R}^{n\times n_{\ell}},\quad\ell\in\{1,\ldots,L\}.$
Then, the conjugate kernel matrix is defined as the Gram matrix of features produced by the final layer of a network, i.e., $X_{L}X_{L}^{\top}\in\mathbb{R}^{n\times n}$ [fan_ck]. The conjugate kernel is central to the analysis of random features models. This connection stems from the fact that the output of a neural network is linear in those derived features. Thus, the conjugate kernel characterizes the training and test error of this linear model. In the case of shallow random features models ($L=1$), the conjugate kernel is equivalent to the Gram matrix of the random features themselves. Indeed, in the setting of Theorem 2.3, the conjugate kernel is given by $AA^{\top}$.
Numerous studies have employed random matrix theory to investigate the conjugate kernel in the proportional regime. For shallow random features models with isotropic data and weight matrices, works such as [pennington_ck, benigni_moments-method] use the moment method to establish deterministic equivalents for the conjugate kernel, while [piccolo2021analysis] extends these results to include an additive bias. More generally, [chouard2022quantitative, louart_random-matrix-approach] analyze the conjugate kernel of random features models with deterministic data, relying on concentration of measure and leave-one-out techniques. Moving beyond the shallow setting, [fan_ck] introduces a notion of approximate pairwise orthogonality for the data that propagates through network layers, enabling the analysis of deep random features models (see Section 2.4.4). Their framework also covers isotropic data as a special case, with the difference that the data are conditioned upon. Beyond bulk spectral properties, motivated by results such as [ba2022highdimensional, cui2024asymptotics], where a single gradient step at random initialization is shown to be equivalent to a spiked random features model, recent works have investigated outlier eigenvalues of the conjugate kernel [benigni2022largest, wang2024nonlinear]. This structure in random features model turns out to have important implications, notably in understanding the mechanisms of feature learning.
The proof of Theorem 2.3 recovers a deterministic equivalent for the conjugate kernel $AA^{\top}$, consistent with the findings of [chouard2022quantitative, louart_random-matrix-approach], employing a significantly different approach. In particular, replacing the ridge parameter by a spectral parameter in Theorem 2.3 and using the Stieltjes inversion lemma, we can recover the density of the empirical spectral distribution of the conjugate kernel.
@@ 3.2 Dyson Equation and Linearizations
The Dyson equation is a self-consistent equation that has proven to be a valuable tool in random matrix theory. For an excellent introduction to this subject, we recommend [erdos2019matrix]. It has been particularly crucial in establishing local laws for Wigner-type matrices, which are self-adjoint random matrices with independent centered entries and a prescribed variance profile [Ajanki_2017_vector, Ajanki_2019_vector1, Ajanki_2019_vector2]. In this setting, the Dyson equation collapses to a scalar one when the variance profile is stochastic (corresponding to the generalized Wigner case) or becomes vector-valued in the more general setting, in which case it is referred to as the quadratic vector equation in [Ajanki_2017_vector]. These works assume a flatness condition on the variance profile, which roughly enforces a mean-field scaling of the variances, together with boundedness of the vector solution. The latter property is easily verified in the spectral bulk but requires more delicate analysis near the spectral edges. Under these assumptions, the vector Dyson equation has been studied in great detail, and for instance it has been established that the associated self-consistent density of states is uniformly $1/3$-Hölder continuous with respect to the Lebesgue measure and supported on finitely many intervals, called bands [alt_band]. Overall, this line of work provides a comprehensive analysis of the vector Dyson equation and its implications for random matrices, notably optimal local laws in both the bulk and at the spectral edges. In general, such local laws are proved by combining a stability analysis of the vector Dyson map with some fluctuation-averaging techniques to control random fluctuations of the resolvent around its deterministic equivalent. While the flatness assumption covers a broad class of variance profiles, it excludes important structured examples such as band matrices or Hermite dilations, which require different methods.
Going beyond Wigner-type matrices, the Dyson equation has also been employed to study random matrices with correlated entries [Alt_2020, ERD_S_2019, ajanki_stability_2019]. In this line of work, one considers real symmetric or complex Hermitian random matrices of the form $H=A+W$, where $A$ is a deterministic bare matrix and $W$ is a random perturbation with centered entries that may exhibit correlations subject to suitable moment bounds. In contrast to the independent-entry case, these correlations are assumed to be local, in the sense that they decay with a metric distance between indices. A flatness condition is also imposed on the covariance structure, ensuring a mean-field scaling of the correlations analogous to the variance flatness in the vector Dyson setting. Under these assumptions, the Dyson equation no longer reduces to a vector equation but remains genuinely matrix-valued, giving rise to the matrix Dyson equation (MDE). The stability for the MDE was developed in [ajanki_stability_2019], providing a foundation for further analysis. Building on this foundation, [ERD_S_2019] proved optimal local laws in the bulk for correlated ensembles by employing a multivariate cumulant expansion to show that the resolvent satisfies the MDE up to a negligible error term. In parallel, [Alt_2020] focused on regular spectral edges, which notably shows the absence of eigenvalues outside the self-consistent support and delocalization of eigenvectors. Although this framework allows for correlated entries, it remains restricted to a decaying correlation structures satisfying the mean-field flatness condition. By contrast, the models considered in this work involve structured correlation patterns that do not necessarily exhibit rapid decay, and our analysis concerns pseudo-resolvents, for which the existing MDE stability techniques cannot be applied directly.
The Dyson equation has also been applied to the study of Kronecker random matrices [alt_Kronecker], which are block-structured random matrices obtained as linear combinations of Kronecker products between deterministic structure matrices and random matrices with independent, centered entries (subject to symmetry constraints). A key feature of this model is that it removes the flatness assumption required in earlier works, imposing only an upper bound on the entries of the variance profile. In this context, the corresponding self-consistent relation is referred to as the Hermitized Dyson equation, which can be viewed as a Kronecker-type extension of the matrix Dyson equation. Using this formulation, [alt_Kronecker] proved that the eigenvalues of such ensembles lie, with high probability, within an $\epsilon$-neighborhood of the self-consistent spectrum defined by the Hermitized Dyson equation. Their stability analysis builds upon the framework developed in [ajanki_stability_2019], suitably adapted to the Kronecker setting. An important distinction from our work is that Kronecker random matrices are linear in the random variables, whereas our models involve polynomial dependence, requiring a linearization step and the analysis of pseudo-resolvents.
Overall, the Dyson equation is a useful tool for deriving deterministic equivalents, but it is limited to certain classes of “Wigner-like” matrices. In order to extend its applicability to a broader class of matrices, we employ a linearization trick. The concept of linearization became particularly important following the influential work of [haagerup_new_2005]. This work essentially demonstrated that to analyze a polynomial expression in matrices, it is sufficient to consider a linear polynomial with matrix coefficients. More precisely, given a polynomial in non-commuting variables, they construct a linearization $L$ such that the original polynomial can be recovered as a Schur complement of the linearization. Because the matrices considered in [haagerup_new_2005] are independent Gaussian Unitary Ensemble matrices, the authors can exploit operator-valued free probability to analyze the asymptotic behavior of the spectrum. Their main result shows that the empirical spectral distribution of polynomials in independent GUE matrices converges, in the large-dimensional limit, to that of the corresponding polynomial in freely independent semicircular variables.
However, a limitation of [haagerup_new_2005] is that their linearization construction does not, in general, preserve the self-adjointness of the original polynomial expression. [Anderson_2013] addressed this problem, proving that a self-adjoint polynomial expression allows the linearization’s coefficients to be chosen in a way that retains self-adjointness [Anderson_2013]. Using this refined construction, [Anderson_2013] established the convergence of the largest singular value of self-adjoint polynomials in independent square Wigner matrices to the operator norm of the corresponding polynomial in free semicircular variables. Beyond the linearization itself, the analysis employs a version of the Dyson equation for linearizations, referred to there as the Schwinger-Dyson equation. For polynomials in independent Wigner matrices, this equation collapses to a finite-dimensional system whose deterministic solution can be constructed explicitly using tools from free probability. Rather than relying on stability analysis or fluctuation-averaging methods as in the previously mentioned works on the Dyson equation, [Anderson_2013] combined matrix identities with $L^{p}$-type estimates to obtain global convergence results. Furthermore, [BelinschiMaiSpeicher+2017+21+53] combined Anderson’s self-adjoint linearization with operator-valued free probability to prove that the empirical eigenvalue distribution of self-adjoint polynomials in several independent Wigner matrices converges to that of the corresponding polynomial in freely independent semicircular variables. The linearization framework was further generalized in [HELTON20181, HELTON2006105] to cover rational expressions in non-commutative variables, thereby extending the reach of free-probability techniques beyond polynomial settings.
The fact that linearizations are not unique has motivated the development of several refined constructions. In particular, [nemish_local_2020] introduced the concept of a minimal linearization, providing a systematic way to construct a self-adjoint linearization of smallest possible size for a given polynomial. This paper considers self-adjoint polynomials in several independent square random matrices with centered entries of identical variance. Within the Dyson equation for linearizations framework, the authors establish a local law for such models. A key technical insight is the identification of a nilpotent structure within the minimal linearization, which allows them to control norm of the pseudo-resolvent. Instead of the $L^{p}$-type estimates used in [Anderson_2013], their proof relies on establishing stability of the DEL itself, combined with large-deviation and fluctuation-averaging estimates adapted from earlier Dyson-equation analyses. As in [Anderson_2013], the DEL reduces in their setting to a finite-dimensional system of equations, since the matrices have independent entries with equal variance. The existence of the deterministic solution can therefore be stated explicitely using tools from free-probability. Analogously to our own approach, their argument involves working with a regularized DEL, where a small imaginary part is added to stabilize the inverse. They derive uniform bounds in this regularized setting before removing the regularization.
The local law proved in [nemish_local_2020] extends the earlier work [anderson_anticommutator], which focuses on a specific polynomial, namely the anticommutator $XY+YX$ of two independent Wigner matrices. In that setting, the author formulates and analyzes a DEL tailored to the specific self-adjoint linearization of the anticommutator, proving existence and stability of the fixed-point equation and using these properties to derive a local law for the resolvent. Finally, [fronk2023norm] investigates Hermitian non-commutative quadratic polynomials in multiple independent Wigner matrices, with a particular focus on the spectral edges. They establish a local law near the edges and, building on it, show that the operator norm of such polynomials converges to a deterministic limit. This work can be viewed as a specialization of [nemish_local_2020] to the quadratic case, where the additional structure allows for sharper control near the edges, a regime that is typically more delicate than the bulk.
A key strength of these linearization techniques is that their supporting arguments are often constructive, giving explicit instructions for creating appropriate linearizations. Combined with operator-valued free probability, the linearization trick—referred to as the pencil method in this context—has found successful applications in the study of simple neural networks [mel2022anisotropic, Adlam2019ARM, tripledescent, adlam_doubledescent, tripuraneni2021overparameterization].
We conclude this section by briefly highlighting how our setting differs from previous works. First, we consider pseudo-resolvents rather than resolvents, which lack several of the structural and norm-control properties enjoyed by classical resolvents. As a result, the existence and stability properties of the associated Dyson equation are not directly transferable, and new arguments are required to establish them. Second, we allow for structured and possibly rectangular random matrices with correlated entries, whereas prior works on pseudo-resolvents focus on square matrices with independent entries of identical variance. Consequently, the corresponding Dyson equation no longer collapses to a finite-dimensional system, but must instead be treated as an matrix-valued equation.
Third, since the deterministic equivalent is not assumed to satisfy an algebraic equation in free variables, free-probability techniques cannot be used to construct and analyze the deterministic limit. Establishing existence and stability therefore requires a different analytic approach tailored to the pseudo-resolvent framework. Finally, we note that in this work we restrict attention to global laws, whereas [nemish_local_2020, anderson_anticommutator] derive local laws with optimal precision.